# revtex4-2Repair the float

### v1

@@ Abstract
The computational power of symmetry-protected phases of matter can be accessed through local measurements, but what is the most efficient way of doing so? In this work, we show that minimizing operational resources in measurement-based quantum computation on subsystem symmetric resource states amounts to solving a sub-Riemannian geodesic problem between the identity and the target logical unitary. This reveals a geometric structure underlying MBQC and offers a principled route to optimize quantum processing in computational phases.
@@ I Introduction
Quantum computation is an interplay between the continuous and the discrete. For example, the states of any given number of qubits form a continuous set, but their measurement yields discrete outcomes. Universality refers to the continuous groups $\text{SU}(2^{n})$, but fault tolerance only provides discrete sets of protected encoded gates [1].—Fortunately, these gates can be composed to circuits dense in $\text{SU}(2^{n})$, as demonstrated by the Solovay-Kitaev theorem [2].
The efficiency of quantum computation is typically based on discrete measures, such as qubit and gate counts, and circuit depth. Yet, a continuous notion also exists [3, 4], due to Nielsen et al., and is called geometric complexity. They consider the optimization of execution time for unitary gates, as generated by time-dependent Hamiltonians. Hard-to-implement terms in the Hamiltonian—for example terms that correspond to many-body interactions—are eliminated as generators of gates. The resulting control theory points to sub-Riemannian geometry.
@@
In 2D, the physical orders known to be conducive to MBQC are subsystem SPT and subsystem SET physical order; and the corresponding computational order parameters $\sigma$ witness those orders. Computational universality can arise in both scenarios; see [11, 14, 15, 16] and [13], respectively.
@@
The remainder of this paper is organized as follows. In Sec. II) we review from [12] the noise model for logical operations in computational phases of quantum matter, and define sub-Riemannian geodesics. With these notions, in Sec. III we formulate our main theorem (Theorem 1) that extends the error bound for MBQC-implemented gates to arbitrary unitaries. In Sec. IV we prove Theorem 1, and illustrate it through the example of $\text{SU}(2)$. Sec. V is our Conclusion and Outlook.
@@ II Preliminaries
In computational phases of quantum matter, non-trivial gates are implemented via measurements in a rotated basis away from the symmetry respecting axis. In case of an ideal resource state—a central state in computational phase considered—each of these non-trivial measurements $M_{i}$, $i=1,\dots N$ implements a local target unitary $U_{i}(\alpha_{i})=\text{e}^{\text{i}\alpha_{i}g_{i}}$ with $g_{i}\subset G_{i}$ the set of locally executable generators and $\alpha_{i}$ the rotation angle. The set $\mathcal{G}=\bigcup_{i=1}^{N}G_{i}$ of executable generators is given by all locally executable generators. The target unitary is implemented as a sequence of rotations $U_{i}=\text{e}^{\text{i}\alpha_{i}g_{i}}$ with angles $\alpha_{i}$, generators $g_{i}$ and resulting unitary gate $U=\prod_{i=1}^{N}U_{i}$.
Away from the perfect resource states, deeper into the phase, the effect of the local measurements remains a product of quantum channels $C=\prod_{i=1}^{N}C_{i}$, under the additional assumption that the symmetry-breaking measurements are far enough apart (longer than the correlation length $\zeta$ of the resource state). The individual logical channels then take the form (see [12, p. 23])
@@
If the Lie algebra generated by $\mathcal{G}$ is $\mathfrak{su}(2^{n})$, we know that we can execute all unitaries $U\in\text{SU}(2^{n})$. This is due to Chows Theorem [19], a result from sub-Riemannian geometry applied to the special unitary group $\text{SU}(2^{n})$. If $\mathcal{G}$ is not bracket-generating, it generates a Lie subalgebra. The resulting circuits then form some closed Lie subgroup of $\text{SU}(2^{n})$.
@@
$\dot{C}(t)=C(t)c(t),\;c(t)\in\text{span}(\text{i}\mathcal{G}).$
Therein, the control Hamiltonian $h(t)=-\text{i}c(t)$ is a linear combination $h(t)=\sum_{g\in\mathcal{G}}c_{g}(t)g$ of executable generators. The functions $c_{g}$ are called controls.
Given $U\in G$ a target unitary, then a connecting horizontal curve additionally satisfies the boundary conditions
C(0)=1,C(1)=U.C(0)=1,\;C(1)=U.
@@
$L(C)=\int_{0}^{1}|c(t)|\,\text{d}t=\int_{0}^{1}\sqrt{\sum_{g\in\mathcal{G}}c_{g}^{2}(t)}\;\,\text{d}t$
@@
For a minimizing curve $C$, the sub-Riemannian arclength $L(C)$ is the sub-Riemannian (Carnot-Caratheodory) distance $\text{d}_{CC}(1,U)$ of $1$ and $U$.
@@ III Result
In the following theorem, we generalize the result of [20, Theorem 1]. Whereas the existing theorem refers only to unitaries generated by a fixed Pauli operator, the present result refers to all unitaries reachable.
@@
Consider a resource state with a translation invariant computational order parameter $\sigma$ and executable generators $\mathcal{G}$. Let $A$ the Lie hull of the executable generators $\mathcal{G}$ and $G$ the corresponding closed Lie subgroup. Let $N$ the number of sites with non-trivial measurement bases; i.e., the splitting number. Further assume that the distance between symmetry breaking measurements is much larger than the correlation length, such that their effects are uncorrelated.
@@
$\epsilon\leq\frac{1}{N}\left(\frac{1}{\sigma^{2}}-1\right)\,\text{d}^{2}_{CC}(e,U)+\mathcal{O}(N^{-2}).$ (3)
Therein, $\text{d}_{CC}(e,U)$ is the sub-Riemannian distance (Def. 1) of $U$ to the identity $e$.
@@ IV Methods
Therefore we can expand the super operators $\text{e}^{\text{i}\alpha g_{i}}]$ in orders of $\alpha=\alpha_{i},\beta_{i}$, to get
$1+\text{i}\alpha[g_{i},\cdot]+\frac{1}{2}\alpha^{2}[g_{i},\cdot]^{2}+\mathcal{O}(\alpha^{3}).$
@@
The contribution of the rotation sequence in the total error $\sum_{i=1}^{N}\alpha_{i}^{2}$ can be seen as a discrete version of the integral $\int_{0}^{1}\sum_{g}c_{g}(t)^{2}\,\text{d}t$, which resembles the sub-Riemannian energy (2).
@@
$U=\prod_{i=1}^{N}\text{e}^{\text{i}\alpha_{i}g_{i}}$
is the end-point of a piecewise-one-parameter curve $C:[0,N]\to\text{SU}(2^{n})$. So for every rotation, we get a piece of the curve $t\in(i-1,i]$:
$C(t)=\prod_{1<j<i}\text{e}^{\text{i}\alpha_{j}g_{j}}\text{e}^{\text{i}(t-i+1)\alpha_{i}g_{i}}.$
Moreover the piecewise-constant controls
$c(t)=\text{i}\,\alpha_{i}g_{i},\quad t\in(i-1,i]$
@@
$\lim_{N\to\infty}N\cdot\epsilon(N)=\left(\frac{1}{\sigma^{2}}-1\right)\int_{0}^{1}|c(t)|^{2}\text{d}t.$
@@
$T^{2}=\left(\int_{0}^{T}|c(t)|\,\text{d}t\right)^{2}=T\int_{0}^{T}|c(t)|^{2}\,\text{d}t$
Here, we applied the Cauchy-Schwarz inequality with $g(t)=1$. Moreover, since $|c(t)|$ is constant, equality holds. Therefore minimizing the arclength (the square $T^{2}$) is equivalent to minimizing the leading $\frac{1}{N}$ coefficient of the error. ∎
Since the controls cannot be applied continuosly, we approximate the geodesics for a good error scaling using a Trotter-like product formula for the pieces of the geodesic. These need to vanish faster in terms of $N$, so they can be neglected compared to the implementation error.
There are som choices for higher order Trotter like product formulae, which also include the commutator and higher order error [22, 5]. One can also use [23] for the time-ordered exponential, or other splitting techniques [24].
We adapt a scheme from [25] to get a local approximation of third order. This results in a global approximation of second order, which is sufficient assuming that the error in the implementation is dominant, of first order.
@@
Compare [25, 5], let $c\left(t\right)=\sum_{g\in\mathcal{G}}c_{g}\left(t\right)g$ be a time-dependent Hamiltonian, and label the generators $g_{i}\in\mathcal{G}$ with $i=1,\dots,m$. Then its first order Lie-Trotter-Suzuki formula is
$\begin{split}&U_{1}\left(t+\Delta,t\right)\equiv\\ &\prod_{i=1}^{m}\exp c_{i}\left(t+\frac{\Delta}{2}\right)\frac{\Delta}{2}\prod_{i=m}^{1}\exp c_{i}\left(t+\frac{\Delta}{2}\right)\frac{\Delta}{2}.\end{split}$ (5)
In figure 2 we compare the resulting curve for the rotation sequence generated by the Lie-Trotter-Suzuki formula to the original geodesic.
@@
Figure 2: The sub-Riemannian geodesic (Example 1) for a rotation in the $y$ direction (Purple). For that curve we compare the Lie-Trotter-Suzuki product formula (Green), which is a curve with piecewise constant controls. Below, the control functions $c_{z}(t)$ (purple) and $c_{x}(t)$ (green) are depicted.
@@
Compare [25, 9], let $U=\mathscr{T}\exp\left[\int\,\text{d}t\,c\left(t\right)\right]$ be the unitary generated by the continuous time-dependent Hamiltonian $c\left(t\right)$. The first-order Lie-Trotter-Suzuki product formula (LABEL:eq:firstorderLieTrotterSuzuki) has a local error of third order. In other words $\left\|U-U_{1}\right\|\in\mathcal{O}\left(\Delta^{3}\right)$.
@@
We start Taylor expanding both parts, the exact rotation using [25, Lemma 1, Lemma 7]
@@
We summarize this, stating how a unitary $U$ is implemented on a ressource state with $N\cdot\Delta$ qubits.
@@
Let $U\in\mathcal{SU}(2^{n})$ and $C(t)$ a minimizing curve with controls $c(t)$. Split the geodesic into $M$ pieces with $M=\lfloor\frac{N}{(2m-1)}\rfloor$. Then we get a rotation sequence containing less than $N$ individual rotations with angles of order $\mathcal{O}(N^{-1})$, according to (Eq. LABEL:eq:firstorderLieTrotterSuzuki).
@@
All the generators in the rotation sequence are executable $c_{g}(t_{i})g$, because we use a horizontal curve $c(t)\in\text{i}\;\text{span}\,\mathcal{G}$.
@@
$\left\|\prod_{i}\text{e}^{\text{i}\alpha_{i}g_{i}}-U\right\|\leq\sum_{i}\mathcal{O}(N^{-3})\in\mathcal{O}(N^{-2}).$
@@
$\sum_{i}|\alpha_{i}|^{2}\leq\frac{1}{N}\int_{0}^{1}|c(t)|^{2}\text{d}t=\frac{1}{N}\,\text{d}^{2}_{CC}(e,U).$
@@ 1 Scaling
Since computing the sub-Riemannian distance relies on finding shortest geodesics, the computation of the distance and some optimal controls is hard. The advantage of the following $\infty$-distance in comparison to the Carnot-Carathéodory (CC) distance is, that it can be calculated more easily for arbitrary unitaries. Basically only the degree of the used generators is of importance. Let $X_{i}$ an orthonormal basis of the Lie algebra. The degree $\deg X_{i}$ of $X_{i}$, i.e. the number of commutators needed to generate $X_{i}$ from $\mathcal{G}$.
@@
$C_{1}\,\text{d}_{\infty}(1,U)\leq\,\text{d}_{CC}(1,U)\leq C_{2}\,\text{d}_{\infty}(1,U).$
So for the scaling of the sub-Riemannian distance we only need to know the scaling of the $\infty$ metric. The inequality for the scaling $\text{d}_{\text{CC}}\leq C|y_{i}|^{\frac{1}{N}}$ can be shown using the BCH formula and is sometimes stated as a consequence or a by-product of Chows theorem for reachability [19, p. 28].
@@
Another way to think about upper bounds is by constructing other curves that implement unitaries with the same restriction. One could implement arbitrary unitaries using a generalization of Euler angles. In [28], they find rotation sequences for exponentials of single Paulis with length $t+(\text{deg}X_{i}-1)\pi$. They also use that in MBQC $\frac{\pi}{2}$ rotations by Pauli operators can be done by classical post-processing.
$\text{e}^{\text{i}\frac{\pi}{2}P_{1}}\dots\text{e}^{\text{i}\frac{\pi}{2}P_{L-1}}\text{e}^{\text{i}tP_{L}}\text{e}^{-\text{i}\frac{\pi}{2}P_{L-1}}\dots\text{e}^{-\text{i}\frac{\pi}{2}P_{1}}.$
Implementing that sequence gives long sequences of a single type of rotation and introduces an additive overhead to implement that sequence for every single Pauli generator appearing in the unitary. That means, that especially for unitaries close to the identity, the scaling argues in favor of the sub-Riemannian minimizing curves. In figure 4 we see as an example that even for a single generator the Euler angles are inefficient compared to using geodesic. With regards to the executable generators we come to the same conclusion as [28]: Optimal generation of the Lie algebra using commutators is benefitial for optimal generation of unitaries using only executable generators. This is due to the scaling of the length of both measures with the degree of the generators.

### v2

# Sub-Riemannian geometry of measurement based quantum computation

@@ Abstract
The computational power of quantum phases of matter with symmetry can be accessed through local measurements, but what is the most efficient way of doing so? In this work, we show that minimizing operational resources in measurement-based quantum computation (MBQC) on subsystem symmetric resource states amounts to solving a sub-Riemannian geodesic problem between the identity and the target logical unitary. This reveals a geometric structure underlying MBQC and offers a principled route to optimize quantum processing in computational phases.
@@ I Introduction
Quantum computation is an interplay between the continuous and the discrete. For example, the states of any given number of qubits form a continuous set, but their measurement yields discrete outcomes. Universality refers to the continuous groups $\mathrm{SU}(2^{n})$, but fault tolerance only provides discrete sets of protected encoded gates [14]. Fortunately, these gates can be composed to circuits dense in $\mathrm{SU}(2^{n})$, as demonstrated by the Solovay-Kitaev theorem [6].
The efficiency of quantum computation is typically based on discrete measures, such as qubit and gate counts, and circuit depth. Yet, a continuous notion also exists [26, 27], due to Nielsen et al., and is called geometric complexity. They consider the optimization of execution time for unitary gates, as generated by time-dependent Hamiltonians. Hard-to-implement terms in the Hamiltonian—for example terms that correspond to many-body interactions—are eliminated as generators of gates. The resulting control theory points to sub-Riemannian geometry. In fact, there are numerous examples in optimal quantum control theory where sub-Riemannian geometries are helpful. Most prominent is the implementation of time optimal quantum gates in specific physical scenarios [20, 21].
@@
In 2D, the physical orders known to be conducive to MBQC are subsystem symmetry-protected topological and subsystem symmetry-enriched topological physical order; and the corresponding computational order parameters $\sigma$ witness those orders. Computational universality can arise in both scenarios; see [29, 12, 36, 11] and [18], respectively.
@@
The remainder of this paper is organized as follows. In Sec. II) we review from [31] the noise model for logical operations in computational phases of quantum matter, and define sub-Riemannian geodesics. With these notions, in Sec. III we formulate our main theorem (Theorem 1) that extends the error bound for MBQC-implemented gates to arbitrary unitaries. In Sec. IV we prove Theorem 1, and illustrate it through the example of $\mathrm{SU}(2)$. Sec. V is our Conclusion and Outlook.
@@ II Preliminaries
In computational phases of quantum matter, non-trivial gates are implemented via measurements in a rotated basis away from the symmetry respecting axis. In case of an ideal resource state—a central state in computational phase considered—each of these non-trivial measurements $M_{i}$, $i=1,\dots N$ implements a local target unitary $U_{i}(\alpha_{i})=\mathrm{e}^{\mathrm{i}\alpha_{i}g_{i}}$ with $g_{i}\subset G_{i}$ the set of locally executable generators and $\alpha_{i}$ the rotation angle. The set $\mathcal{G}=\bigcup_{i=1}^{N}G_{i}$ of executable generators is given by all locally executable generators. The target unitary is implemented as a sequence of rotations $U_{i}=\mathrm{e}^{\mathrm{i}\alpha_{i}g_{i}}$ with angles $\alpha_{i}$, generators $g_{i}$ and resulting unitary gate $U=\prod_{i=1}^{N}U_{i}$.
Away from the perfect resource states, deeper into the phase, the effect of the local measurements remains a product of quantum channels $C=\prod_{i=1}^{N}C_{i}$, under the additional assumption that the symmetry-breaking measurements are far enough apart (longer than the correlation length $\zeta$ of the resource state). The individual logical channels then take the form (see [31, p. 23])
@@
If the Lie algebra generated by $\mathcal{G}$ is $\mathfrak{su}(2^{n})$, we know that we can execute all unitaries $U\in\mathrm{SU}(2^{n})$. This is due to Chows Theorem [32], a result from sub-Riemannian geometry applied to the special unitary group $\mathrm{SU}(2^{n})$. If $\mathcal{G}$ is not bracket-generating, it generates a Lie subalgebra. The resulting circuits then form some closed Lie subgroup of $\mathrm{SU}(2^{n})$.
@@
$\dot{C}(t)=C(t)c(t),\;c(t)\in\mathrm{span}(\mathrm{i}\mathcal{G}).$
Therein, the control $c(t)$ is a linear combination
$c(t)=\mathrm{i}\sum_{g\in\mathcal{G}}c_{g}(t)g$
of executable generators. The functions $c_{g}$ are called control functions for the generator $g\in\mathcal{G}$ and form a vector $\vec{c}(t)=(c_{g})_{g\in\mathcal{G}}$. Given $U\in G$ a target unitary, then a connecting horizontal curve additionally satisfies the boundary conditions
C(0)=e,C(1)=U.C(0)=e,\;C(1)=U.
@@
$L(C)=\int_{0}^{1}|\vec{c}(t)|\,\text{d}t=\int_{0}^{1}\sqrt{\sum_{g\in\mathcal{G}}c_{g}^{2}(t)}\;\,\text{d}t$
@@
For a minimizing curve $C$, the sub-Riemannian arclength $L(C)$ is the sub-Riemannian (Carnot-Caratheodory) distance $\text{d}_{\mathrm{CC}}(e,U)$ of $e$ and $U$.
@@ III Result
In the following theorem, we generalize the result of [1, Theorem 1]. Whereas the existing theorem refers only to unitaries generated by a fixed Pauli operator, the present result refers to all unitaries reachable.
@@
Consider a resource state with a translation invariant computational order parameter $\sigma$ and executable generators $\mathcal{G}$. Let $A$ be the Lie hull of the executable generators $\mathcal{G}$ and $G$ be the corresponding closed Lie subgroup. Let $N$ be the number of sites with non-trivial measurement bases; i.e., the splitting number. Further assume that the distance between symmetry breaking measurements is much larger than the correlation length, such that their effects are uncorrelated.
@@
$\epsilon\leq\frac{1}{N}\left(\frac{1}{\sigma^{2}}-1\right)\,\text{d}^{2}_{\mathrm{CC}}(e,U)+\mathcal{O}(N^{-2}).$ (3)
Therein, $\text{d}_{\mathrm{CC}}(e,U)$ is the sub-Riemannian distance (Def. 1) of $U$ to the identity $e$.
@@ IV Methods
Therefore we can expand the super operators $\mathrm{e}^{\mathrm{i}\alpha g_{i}}]$ in orders of $\alpha=\alpha_{i},\beta_{i}$, to get
$1+\mathrm{i}\alpha[g_{i},\cdot]+\frac{1}{2}\alpha^{2}[g_{i},\cdot]^{2}+\mathcal{O}(\alpha^{3}).$
@@
The contribution of the rotation sequence in the total error $\sum_{i=1}^{N}\alpha_{i}^{2}$ can be seen as a discrete version of the integral $\int_{0}^{1}\sum_{g}c_{g}(t)^{2}\,\text{d}t$, which resembles the sub-Riemannian energy (Lemma 2).
@@
$U=\prod_{i=1}^{N}\mathrm{e}^{\mathrm{i}\alpha_{i}g_{i}}$
is the end-point of a piecewise-one-parameter curve $C:[0,N]\to\mathrm{SU}(2^{n})$. So for every rotation, we get a piece of the curve $t\in(i-1,i]$:
$C(t)=\prod_{1<j<i}\mathrm{e}^{\mathrm{i}\alpha_{j}g_{j}}\mathrm{e}^{\mathrm{i}(t-i+1)\alpha_{i}g_{i}}.$
Moreover the piecewise-constant control
$c(t)=\mathrm{i}\,\alpha_{i}g_{i},\quad t\in(i-1,i]$
@@
$\lim_{N\to\infty}N\cdot\epsilon(N)=\left(\frac{1}{\sigma^{2}}-1\right)\int_{0}^{1}|c(t)|^{2}\mathrm{d}t.$
@@
$T^{2}=\left(\int_{0}^{T}|c(t)|\,\text{d}t\right)^{2}=T\int_{0}^{T}|c(t)|^{2}\,\text{d}t.$
Here, we applied the Cauchy-Schwarz inequality with $g(t)=1$. Moreover, since $|c(t)|$ is constant, equality holds. Therefore minimizing the arclength (the square $T^{2}$) is equivalent to minimizing the leading $\frac{1}{N}$ coefficient of the error bound. ∎
Since the control cannot be applied continuosly, we approximate the geodesics for a good error scaling using a Trotter-like product formula for the pieces of the geodesic. These need to vanish faster in terms of $N$, so they can be neglected compared to the implementation error.
There are some choices for higher-order Trotter like product formulae, which also include the commutator and higher-order error [8, 5]. One can also use [37] for the time-ordered exponential, or other splitting techniques [7].
We adapt a scheme from [38] to get a local approximation of third order. This results in a global approximation error of second order, which is sufficient, because the error in the implementation is of first order, so it is dominant. We recall from [38], Page 5 the following definition.
@@
Let $c$ be a time-dependent control
$c\left(t\right)=\mathrm{i}\sum_{g\in\mathcal{G}}c_{g}\left(t\right)g.$
We label the generators $g_{i}\in\mathcal{G}$ with $i=1,\dots,m$. Then its first order Lie-Trotter-Suzuki formula is
[…]
In figure 2 we compare the resulting curve for the rotation sequence generated by the Lie-Trotter-Suzuki formula to the original geodesic. The local approximation order (Lemma 3) is also given in [38], Page 9.
@@
Figure 2: The sub-Riemannian geodesic (Example 1) for a rotation in the $y$ direction (Purple). For that curve we compare the Lie-Trotter-Suzuki product formula (Green), which is a curve with piecewise constant control. To the right, the control functions $c_{z}(t)$ (purple) and $c_{x}(t)$ (green) are depicted.
@@
Let $U=\mathscr{T}\exp\left[\int\,\text{d}t\,c\left(t\right)\right]$ be the unitary generated by the continuous time-dependent control $c\left(t\right)$. The first-order Lie-Trotter-Suzuki product formula (LABEL:eq:firstorderLieTrotterSuzuki) has a local error of third order. In other words $\left\|U-U_{1}\right\|\in\mathcal{O}\left(\Delta^{3}\right)$.
@@
We start Taylor expanding both parts, the exact rotation using [38, Lemma 1, Lemma 7]
@@
We summarize this, stating how a unitary $U$ is implemented on a resource state with $N\cdot\Delta$ qubits.
@@
Let $U\in\mathrm{SU}(2^{n})$ the target unitary and let $C(t)$ be a minimizing geodesic with control $c(t)$ such that $C(T)=U$. Split the geodesic into $M$ pieces with $M=\lfloor\frac{N}{(2m-1)}\rfloor$. Then we get a rotation sequence containing less than $N$ individual rotations with angles of order $\mathcal{O}(N^{-1})$, according to (Eq. LABEL:eq:firstorderLieTrotterSuzuki).
@@
All the generators in the rotation sequence are executable $c_{g}(t_{i})g$, because we use a horizontal curve $c(t)\in\mathrm{i}\;\mathrm{span}\,\mathcal{G}$.
@@
$\left\|\prod_{i}\mathrm{e}^{\mathrm{i}\alpha_{i}g_{i}}-U\right\|\leq\sum_{i}\mathcal{O}(N^{-3})\in\mathcal{O}(N^{-2}).$
@@
$\sum_{i}|\alpha_{i}|^{2}\leq\frac{1}{N}\int_{0}^{1}|c(t)|^{2}\mathrm{d}t=\frac{1}{N}\,\text{d}^{2}_{\mathrm{CC}}(e,U).$
@@ 1 Scaling
Since computing the sub-Riemannian distance relies on finding shortest geodesics, the computation of the distance and some optimal control is hard. The advantage of the following $\infty$-distance in comparison to the Carnot-Carathéodory (CC) distance is, that it can be calculated more easily for arbitrary unitaries. Basically only the degree of the used generators is of importance. Let $X_{i}$ an orthonormal basis of the Lie algebra. The degree $\deg X_{i}$ of $X_{i}$, i.e. the number of commutators needed to generate $X_{i}$ from $\mathcal{G}$.
@@
$C_{1}\,\text{d}_{\infty}(e,U)\leq\,\text{d}_{\mathrm{CC}}(e,U)\leq C_{2}\,\text{d}_{\infty}(e,U).$
So for the scaling of the sub-Riemannian distance we only need to know the scaling of the $\infty$ metric. The inequality for the scaling $\text{d}_{\mathrm{CC}}\leq C|y_{i}|^{\frac{1}{N}}$ can be shown using the BCH formula and is sometimes stated as a consequence or a by-product of Chows theorem for reachability [32, p. 28].
@@
Another way to think about upper bounds is by constructing other curves that implement unitaries with the same restriction. One could implement arbitrary unitaries using a generalization of Euler angles. In [34], they find rotation sequences for exponentials of single Paulis with length $t+(\mathrm{deg}X_{i}-1)\pi$. They also use that in MBQC $\frac{\pi}{2}$ rotations by Pauli operators can be done by classical post-processing.
$\mathrm{e}^{\mathrm{i}\frac{\pi}{2}P_{1}}\dots\mathrm{e}^{\mathrm{i}\frac{\pi}{2}P_{L-1}}\mathrm{e}^{\mathrm{i}tP_{L}}\mathrm{e}^{-\mathrm{i}\frac{\pi}{2}P_{L-1}}\dots\mathrm{e}^{-\mathrm{i}\frac{\pi}{2}P_{1}}.$
Implementing that sequence gives long sequences of a single type of rotation and introduces an additive overhead to implement that sequence for every single Pauli generator appearing in the unitary. That means, that especially for unitaries close to the identity, the scaling argues in favor of the sub-Riemannian minimizing curves. In figure 4 we see as an example that even for a single generator the Euler angles are inefficient compared to using geodesic. With regards to the executable generators we come to the same conclusion as [34]: Optimal generation of the Lie algebra using commutators is beneficial for optimal generation of unitaries using only executable generators. This is due to the scaling of the length of both measures with the degree of the generators.